OpenStudy (anonymous):

make a table of values and a hand drawn graph for the function y= 1.5^(x)

OpenStudy (anonymous):

y(x)=1.5^x Y(-2)=1.5^(-2) ...(-2,0.444) Y(-1)=1.5^(-1) ...(-1,0.666) Y(0)=1.5^(-0) ...(0,1) Y(1)=1.5^(1) ...(1,1.5) Y(2)=1.5^(2) ...(2,2.25) and so on

OpenStudy (anonymous):

i dont get how u did that

OpenStudy (anonymous):

we chose a few numbers of our own choice (these are -2,-1,0,1,2) and substitute into the given equation we have (replacing x)

OpenStudy (anonymous):

x is an exponent tho

OpenStudy (anonymous):

yes, no problem, we replaced x with a specific number to evaluate the function value at that number

OpenStudy (anonymous):

how do you make a table of values thouhh

OpenStudy (anonymous):